a father 34 years older than his son tony.
their total ages is 54 years
how old will to tony be in 25 years time

Answer by stanbon(75887)

a father 34 years older than his son tony.
their total ages is 54 years
-----
Equations:
f + t = 54 years
f = t + 34
------
how old will to tony be in 25 years time
---
Substitute for "f" and solve for "t":
t+34 + t = 54
2t = 20
t = 10 (Tony's age is now 10)
In 25 years Tony will be 35.
==============
Cheers,
Stan H.